# AI-Empowered Integrated Sensing and Communications

### v1

@@ II-A Communications
Communication systems involve the transmission of signals, such as voice, video, and data from a sender to a receiver through various physical media, including cables, optical fibers, or wireless channels. These systems rely on a series of processes, such as source and channel encoding, modulation, multiplexing, beamforming and waveform design to prepare the information for efficient transmission over the selected medium. At the receiver’s end, the reverse operations are applied to reconstruct the original signal.
@@
5G networks are designed to support diverse applications by providing enhanced performance metrics. Building upon the advancements of 5G, 6G technology aims to significantly enhance key performance indicators. Projected targets for 6G include a peak data rate of 1 terabit per second, an experienced data rate of 1 gigabit per second, and a peak spectral efficiency of 60 bits per second per Hertz (bits/s/Hz) [35]. Additionally, 6G aspires to achieve user plane latency as low as 100 microseconds ($\mu$s) and reliability levels reaching $10^{-7}$. These ambitious goals are designed to support applications such as immersive communications, hyper-reliable low-latency services, autonomous driving, and the extensive connectivity required for the IoT.
@@ II-A1 Communications Performance Metrics
The main figure of merit in digital communication is the energy per bit to noise power spectral density ratio, ${E_{b}}/{N_{0}}$, where $E_{b}$ denotes the energy per bit and $N_{0}$ corresponds to the noise power spectral density. It is related to the SNR by ${E_{b}}/{N_{0}}=\text{SNR}\cdot{W}/{R_{b}}$, where $W$ denotes the system bandwidth, and $R_{b}$ is the bit rate [36]. The ratio ${R_{b}}/{W}$ is known as bandwidth efficiency, or spectral efficiency, and is measured in bits/s/Hz. With ideal pulse shaping, for $M$-ary digital modulation schemes like QPSK this efficiency is $\mathrm{log}_{2}M$ [37]. Another important performance metric is the bit error rate (BER), which represents the probability of bit errors in the received data. BER is commonly plotted as a function of either $\frac{E_{b}}{N_{0}}$ or SNR.
The achievable data rate refers to the data transmission rate in a channel, considering practical constraints. The capacity of a channel is defined as the highest achievable data rate when the BER goes to zero. For an additive white Gaussian noise (AWGN), Shannon’s formula $\rm C=W\,\log[2]{1+{SNR}}$ defines the capacity in bits/s. There are other important metrics such as throughput, outage probability, error vector magnitude (EVM), signal-to-interference-plus-noise ratio (SINR), and latency, all defined in Table II.
TABLE II: Common Performance Metrics for Communication Systems
Metric Definition Description & Additional Information
BER Probability of bit errors in received data. Lower BER indicates more reliable communication. BER depends on ${E_{b}}/{N_{0}}$ (or SNR), modulation scheme, and coding.
SNR Ratio of signal power to noise power. Higher SNR leads to better detection and lower BER.
SINR Ratio of signal power to the sum of interference and noise power. Higher SINR results in better communication quality, especially in interference-limited networks.
Spectral efficiency Bits transmitted per second per Hz of bandwidth. Higher spectral efficiency allows for more data transmission in limited bandwidth.
Capacity Theoretical upper bound on data transmission rate. Given by Shannon’s formula $\rm C=W\,\log[2]{1+SNR}$, represents the maximum possible data rate with vanishing errors.
Achievable rate Data transmission rate considering practical constraints. Determined by modulation, coding, and BER requirements. It is lower than capacity because of implementation limitations.
Throughput Actual data successfully transmitted and received. Affected by network congestion, retransmissions, and protocol overhead. It is always lower than the achievable data rate.
Outage probability Probability that the received signal falls below a threshold. Important in fading channels to assess communication reliability.
Frame error rate Probability of errors in transmitted frames. Used in packet-based communication to assess reliability.
Latency Delay in data transmission and reception. Low latency is crucial for real-time applications like voice, video, gaming, and autonomous systems.
EVM Deviation between actual and ideal signal constellation points. Lower EVM indicates better modulation accuracy and signal quality.
@@ II-B1 Sensing Performance Metrics
The estimation-theoretic metrics are typically used to characterize the sensing performance. In parameter estimation, the performance of the estimator is commonly evaluated using the MSE metric as $\mathrm{MSE}_{\bm{\theta}}=\mathbb{E}[\|\bm{\theta}-\hat{\bm{\theta}}\|^{2}]$, where $\bm{\theta}$ is a vector of deterministic (true) parameters and $\hat{\bm{\theta}}$ is its estimate. The estimator that minimizes the MSE is termed an MMSE estimator, and it may not be realizable for many practical scenarios. From a practical perspective, we constrain the estimator’s bias to be zero to construct an estimator that minimizes the variance. Such an estimator is termed a minimum variance unbiased estimator and its MSE is just the variance. The variance of any unbiased estimator is lower-bounded by the CRLB; $\mathrm{CRB}=\mathbf{I}^{-1}(\bm{\theta})$, where $\mathbf{I}(\bm{\theta})$ is the Fisher’s information matrix. The CRLB is widely used as a sensing performance metric. The notion of CRB can also be extended to random parameters with a known a-priori distribution. Hence, one can define a posterior CRB for Bayesian estimators as it serves as a lower bound for Bayesian MSE. Moreover, Weiss-Weinstein and Ziv-Zakai bounds are two tighter Bayesian lower bounds that can be derived when the prior distribution of the parameter vector is known.
@@ II-C Radar
Radio detection and ranging, or radar, is a type of sensing technology that uses radio waves to detect and determine the distance, speed, and other characteristics of objects. By transmitting waves towards an object and receiving the reflected signals, radar can operate in an active sensing manner. The time delay and frequency changes (Doppler effect) between transmission and reception provide information about the object’s properties. Radar is widely used in various applications.
{forest}
[…]
Applications [air traffic control] [automotive safety systems] [weather forecasting] [military surveillance] [motion detection & perimeter security ] [fall detection and healthcare ] [security and surveillance] ]
Radars can be classified based on various criteria, including functionality and operational configuration. Common classifications include primary versus secondary, active versus passive, and monostatic, bistatic, or multistatic systems, among others.
Primary radar operates by emitting radio waves towards an object and detecting the echoes reflected back from the object. It is suitable for basic detection but provides limited additional information. Secondary radar also emits radio waves towards an object, typically an aircraft, but requires the aircraft to carry a transponder that responds to the radar signal. This ‘cooperative’ interaction provides detailed information beyond basic detection, including aircraft identification and altitude.
Bistatic, monostatic, and multistatic radars are configurations based on the relationship between the transmitter and receiver elements. Monostatic radars have both transmitter and receiver colocated, simplifying design but limiting stealth detection. Bistatic radars separate these elements, improving stealth detection and accuracy but complicating synchronization. Multistatic radars use multiple transmitters and/or receivers across different locations, enhancing range and resistance to jamming but requiring complex coordination.
A common classification of radars is given in Fig. 2. Active radar systems emit their own electromagnetic waves towards a target and detect the reflections that bounce back from objects. In contrast, passive radar systems do not emit their own waves but rely on external sources of illumination for detection. They detect reflections of signals sent from broadcast stations, communication satellites, other radars, etc. While this method allows for covert operation and reduces the risk of detection by potential adversaries, passive radars face challenges related to signal availability, processing complexity, and operational limitations compared to active radar systems.
Radaractive radar continuous-wave radar unmodulated modulated PMCW radarFMCW radarlinearnonlinearpulsed radarMTI radarDoppler radarpassive radar
Figure 2: A classification of different types of radars.
@@
A continuous wave (CW) radar, in contrast to pulsed radars, transmits continuously, meaning that it listens for target returns simultaneously with transmitting. CW radars can beted or unmodulated. Two prominent examples of modulated CW radars are frequency-modulated continuous wave (FMCW) and phase-modulated continuous wave (PMCW) radars. CW radar offers a longer range and higher resolution than pulsed radar but faces challenges like requiring separate transmitters and receivers, difficulty measuring target distance, and susceptibility to interference and jamming. A detailed comparison of pulsed radars and CW radar systems are given in Table III, while Table IV discusses the advantages and disadvantages of different radar types in general.
TABLE III: Comparison of pulsed radar and Continuous-Wave Radar
Feature Pulsed Radar Continuous-Wave Radar
Operation Transmits short-duration pulses of radio frequency (RF) energy. Simultaneously transmit and receive a continuous wave of RF energy.
Functionality Measures time delay for target range. Measures Doppler shift for target velocity.
Applications Air traffic control, weather monitoring, military surveillance. Speed measurement (police radar), automotive radar (adaptive cruise control).
Antenna usage The same antenna can be used for both transmission and reception (via a duplexer). Typically requires separate transmit and receive antennas.
Advantages Good range resolution, long-range detection capability. High Doppler resolution, simpler design.
Disadvantages Limited Doppler resolution, complexity in pulse compression. Limited range information without additional techniques (e.g., FMCW radar).
TABLE IV: Comparison of Advantages and Disadvantages of Different Radar Types
Type Advantages Disadvantages
Pulsed Radar Simple to implement Good for long-range detection Measures both range and speed Limited range resolution High power consumption Potential dead zones
CW Radar Measures velocity well Lower power consumption Smaller hardware Not good for simultaneous range and velocity measurement Needs a separate transmitter and receiver
FMCW Radar Measures both range and velocity Higher accuracy Low power consumption Complex signal processing Susceptible to interference
Pulse Doppler Radar High precision and resolution in speed and distance measurement Distinguishes stationary and moving targets Operates in multiple modes High power consumption High complexity Vulnerable to deception and jamming
@@ II-C1 Radar Performance Metrics
Typical radar and sensing metrics are summarized in Table V. Detecting targets and estimating parameters are two of the primary functionalities of a radar system. In radar detection, the probability of detection, probability of false-alarm, and resolution are the widely used performance metrics. The range resolution of radar systems is defined as $\Delta R=c/2W$, where $c$ is the speed of light and $W$ is the bandwidth. Similarly, the velocity resolution is defined as $\Delta v=\lambda/(2NT)$, where $\lambda$ is the wavelength, $N$ is the number of coherently processed pulses, and $T$ is the pulse repetition interval. Finally, the half-power angular resolution is defined as $\Delta\theta=0.886\lambda/D_{A}$, where $D_{A}$ is the size of the array’s aperture. The range, velocity/Doppler, and angular accuracy can be defined by using $\Delta R$, $\Delta v$, and $\Delta\theta$.
The range (or time delay) is the only unambiguous parameter that early radar systems could estimate. The range rate (or Doppler) can be estimated by CW radars. The range and Doppler can be estimated simultaneously by pulse–Doppler radars. The angular parameters can be estimated by radar arrays. The trade-offs among range, Doppler, and angular resolutions can be characterized using the multidimensional generalized ambiguity function [38]. Under the narrowband assumption, this function can be factorized into the product of the range-Doppler ambiguity function and the azimuth-elevation angle ambiguity function. The classical ambiguity function for narrowband waveforms was first introduced to understand the resolution properties of time delay and Doppler in [39, 40].
In the seminal works [39, 40], the MSE metric between a known radar waveform $s(t)$ and its frequency-shifted and time-delayed version $s(t-\Delta\tau)e^{-j2\pi\Delta f_{d}t}$ is computed as
[…]
where $\Delta\tau$ is the time delay and $\Delta f_{d}$ is the frequency shift. The only term in (1) that depends on $\Delta\tau$ and $\Delta f_{v}$ has been defined as the classical ambiguity function
[…]
where $s^{*}(t)$ denotes the complex conjugate of $s(t)$.
This two-dimensional ambiguity function can also be interpreted as the matched filter output in the receiver when a delay mismatch $\Delta\tau$ and a Doppler mismatch $\Delta f_{d}$ occur. The value of $\chi(\Delta\tau,\Delta f_{d})$ evaluated at $\Delta\tau=0$ and $\Delta f_{d}=0$ corresponds to the matched filter output without any delay-Doppler mismatch. Thus, the larger the ambiguity function at the origin, the better the time-delay (range) and Doppler resolution. Moreover, the ambiguity function must be much smaller for $\Delta\tau\neq 0$ and $\Delta f_{d}\neq 0$. This is because the correlation between $s(t-\Delta\tau)e^{-j2\pi\Delta f_{d}t}$ and $s(t-\Delta\tau^{\prime})e^{-j2\pi\Delta f_{d}^{\prime}t}$ corresponds to $\chi(\Delta\tau-\Delta\tau^{\prime},\Delta f_{d}-\Delta f_{d}^{\prime})$. There exists a fundamental ambiguity or duality between target resolution in time and frequency. The targets cannot be resolved perfectly in time and frequency simultaneously for a given time–bandwidth product, and this is the reason for its name; the ambiguity function.
TABLE V: Radar Performance Metrics for Radar Systems
Metric Definition Description & Additional Information
Range resolution Ability to distinguish closely spaced targets. Higher resolution allows for better target discrimination.
Doppler resolution Ability to distinguish targets by velocity. Important for detecting moving targets.
Ambiguity function The resolution and ambiguity in range and Doppler. Helps analyze the trade-off between range and Doppler resolution.
Probability of detection ($P_{D}$) Likelihood of correctly detecting a target. Higher $P_{D}$ is desirable for reliable detection.
Probability of false alarm ($P_{\rm FA}$) Likelihood of false target detection. Lower $P_{\rm FA}$ is desirable to reduce false alarms.
ROC $P_{\rm D}$ versus $P_{\rm FA}$ for a given decision threshold. ROC curve evaluates a system’s ability to detect the presence of a signal, target, or object versus making a false alarm.
Clutter rejection Ability to distinguish targets from clutter. Effective clutter rejection improves target detection.
Azimuth/elevation resolution Ability to distinguish targets by angular position. Higher angular resolution allows for better spatial discrimination.
CRLB Theoretical lower bound on the variance of unbiased parameter estimators. Provides a benchmark for the best achievable estimation accuracy (e.g., range, velocity).
SNR Ratio of target signal power to noise power. Higher SNR improves detection performance.
MSE Average MSE between estimated and true parameter values. Measures the accuracy of parameter estimation (e.g., range, velocity). Lower MSE indicates better estimation performance.
@@ II-D1 ISAC Performance Metrics
The performance of ISAC designs can be evaluated by characterizing fundamental trade-offs between sensing and communication metrics [5, 41, 42]. Generally, an ISAC metric combines communication and radar matrices from Tables II and V, along with newly defined notions.
The concept of radar estimation information rate, introduced in [41, 43], parallels the data information rate in communication systems. It treats radar illumination on a target as a form of non-cooperative communication, where parameters such as delay, Doppler, and spatial directions are transmitted. Consequently, the radar channel can be viewed as a non-cooperative communication channel. This information rate is defined as the mutual information between radar and target, derived from the entropy of random estimation parameters and the entropy of estimation uncertainty [41]. Using the data processing inequality, the mutual information between the observation $Y$ and the true parameter $\theta$ is lower-bounded as [34] $I(Y;\theta)\geq\frac{1}{2}\log{\frac{P}{D}},$ where $\theta$ is Gaussian-distributed with variance $P$, and $D$ is the MSE distortion of estimating $\theta$. This bound links the equivalent estimation information rate to the MSE distortion.
@@ III Multi-Objective Optimization: Traditional versus Learning-Based Approaches
ISAC problems are inherently multi-objective optimization problems. Such a problem involves simultaneously optimizing two or more conflicting objectives which can be formulated as
$\underset{}{\text{Minimize}}$ $\quad\bm{f}(x)=[f_{1}(x),f_{2}(x),\ldots,f_{i}(x)]$ (3) Subject to $\quad{g}_{j}(x)=0,\quad j=1,2,\ldots,m$ $\quad{h}_{k}(x)\leq 0,\quad k=1,2,\ldots,n\quad$
in which $f_{i}(x)$ are the objective functions, $g_{j}(x)$ and $h_{k}(x)$ are the equality and inequality constraints, respectively, and $x$ is the decision variable vector.
There are several approaches to solving multi-objective optimization problems. In this section, we will discuss these approaches, classifying them to traditional methods and learning-based techniques.
III-A Traditional Methods
Using the scalarization method [44], we can transform a multi-objective optimization problem into a single-objective problem. This can be done in various ways such as weighted sum method and $\epsilon$-constraint method, as explained below.

### v2

@@ II-A Communications
Communication systems involve the transmission of signals, such as voice, video, and data from a sender to a receiver through various physical media, including cables, optical fibers, or wireless channels. These systems rely on a series of processes, such as source and channel encoding, modulation, multiplexing, beamforming and waveform design to prepare the information for efficient transmission over the selected medium. At the receiver end, reverse operations are applied to reconstruct the original signal.
@@
[…]
@@ II-A1 Communications Performance Metrics
The main figure of merit in digital communication is the energy per bit to noise power spectral density ratio, ${E_{b}}/{N_{0}}$, where $E_{b}$ denotes the energy per bit and $N_{0}$ corresponds to the noise power spectral density. It is related to the SNR by ${E_{b}}/{N_{0}}=\text{SNR}\cdot{W}/{R_{b}}$, where $W$ denotes the system bandwidth, and $R_{b}$ is the bit rate [50]. The ratio ${R_{b}}/{W}$ is known as bandwidth efficiency, or spectral efficiency, and is measured in bits/s/Hz. With ideal pulse shaping, for $M$-ary digital modulation schemes like QPSK this efficiency is $\mathrm{log}_{2}M$ [41]. Other common metrics include the bit error rate (BER), symbol error rate (SER), and block error rate (BLER). BER represents the probability of bit errors in the received data, SER quantifies the probability of symbol misdetection, and BLER measures the probability that an entire message block is received incorrectly. These metrics are typically plotted as functions of $E_{b}/N_{0}$ or SNR.
The achievable data rate refers to the data transmission rate in a channel, considering practical constraints. The capacity of a channel is defined as the highest achievable data rate when the BER goes to zero. For an AWGN channel with bandwidth $W$, Shannon’s formula $\rm C=W\,\mathrm{log}_{2}\left(1+{SNR}\right)$ defines the capacity in bits/s. There are other important metrics such as throughput, outage probability, error vector magnitude (EVM), and latency, all defined in Table III.
@@ II-B1 Sensing Performance Metrics
The estimation-theoretic metrics are typically used to characterize the sensing performance. In parameter estimation, the performance of the estimator is commonly evaluated using the MSE metric as $\mathrm{MSE}_{\bm{\theta}}=\mathbb{E}[\|\bm{\theta}-\hat{\bm{\theta}}\|^{2}]$, where $\bm{\theta}$ is a vector of deterministic (true) parameters and $\hat{\bm{\theta}}$ is its estimate. The estimator that minimizes the MSE is termed an MMSE estimator, and it may not be realizable for many practical scenarios. From a practical perspective, we constrain the estimator’s bias to be zero to construct an estimator that minimizes the variance. Such an estimator is termed a minimum variance unbiased estimator and its MSE is just the variance. The variance of any unbiased estimator is lower-bounded by the CRLB; $\mathrm{CRB}=\mathbf{I}^{-1}(\bm{\theta})$, where $\mathbf{I}(\bm{\theta})$ is the Fisher’s information matrix. The CRLB is widely used as a sensing performance metric. The notion of CRLB can also be extended to random parameters with a known a-priori distribution. Hence, one can define a posterior CRLB for Bayesian estimators as it serves as a lower bound for Bayesian MSE. Moreover, Weiss-Weinstein and Ziv-Zakai bounds are two tighter Bayesian lower bounds that can be derived when the prior distribution of the parameter vector is known.
@@ II-C Radar
Radio detection and ranging, or radar, is a type of sensing technology that uses radio waves to detect and determine the distance, speed, and other characteristics of objects. By transmitting waves towards an object and receiving the reflected signals, radar operates in an active sensing manner. The time delay, frequency shifts (Doppler effect), and the angle of arrival (AoA) of the received signals provide information about the object’s properties. Radar’s applications include air traffic control, automotive safety systems, weather forecasting, military surveillance, motion detection and perimeter security, fall detection and healthcare, as well as general security and surveillance.
Radars are typically classified as either active or passive. Active radar systems emit their own electromagnetic waves towards a target and detect the reflections that bounce back from objects. In contrast, passive radar systems do not emit their own waves but rely on external sources of illumination for detection. They detect reflections of signals sent from broadcast stations, communication satellites, other radars, etc. While this method allows for covert operation and reduces the risk of detection by potential adversaries, passive radars face challenges related to signal availability, processing complexity, and operational limitations compared to active radar systems.
TABLE IV: Radar/Sensing Performance Metrics
Metric Definition and Description References
Range resolution Ability to distinguish closely spaced targets. Higher resolution allows for better target discrimination. [51, 52, 53]
Doppler resolution Ability to distinguish targets by velocity. Important for detecting moving targets. [51, 52]
Ambiguity function The resolution and ambiguity in range and Doppler. Helps analyze the trade-off between range and Doppler resolution. [51]
$P_{D}$ Likelihood of correctly detecting a target. Higher $P_{D}$ is desirable for reliable detection. [54]
$P_{\mathrm{FA}}$ Likelihood of false target detection. Lower $P_{\mathrm{FA}}$ is desirable to reduce false alarms. [54]
Receiver operating characteristic (ROC) $P_{D}$ versus $P_{\mathrm{FA}}$ for a given decision threshold. The ROC curve evaluates a system’s ability to detect the presence of a target versus making a false alarm. [54]
Clutter rejection Ability to distinguish targets from clutter. Effective clutter rejection improves target detection. [51]
Azimuth/elevation resolution Ability to distinguish targets by angular position. Higher angular resolution allows for better spatial discrimination. [53]
CRLB Theoretical lower bound on the variance of unbiased parameter estimators. Provides a benchmark for the best achievable estimation accuracy (e.g., range, velocity). [29]
MSE Average mean squared error between estimated and true parameter values. Lower MSE indicates better estimation performance (e.g., in range or velocity). [29]
@@
A continuous wave (CW) radar, in contrast to pulsed radars, transmits continuously, meaning that it listens for target returns simultaneously with transmitting. CW radars can beted or unmodulated. Two prominent examples of modulated CW radars are frequency-modulated continuous wave (FMCW) and phase-modulated continuous wave (PMCW) radars. CW radar offers a longer range and higher resolution than pulsed radar but faces challenges like requiring separate transmitters and receivers, difficulty measuring target distance, and susceptibility to interference and jamming.
@@ II-C1 Radar Performance Metrics
Radar’s main functionalities are detecting, tracking, imaging, and classifying targets. The key metrics associated with these radar and sensing tasks are summarized in Table IV. In radar detection, the probability of detection ($P_{D}$), probability of false alarm ($P_{\rm FA}$), and resolution are the widely used performance metrics. The range resolution of radar systems is defined as $\Delta R=c/2W$, where $c$ is the speed of light and $W$ is the bandwidth. Similarly, the velocity resolution is defined as $\Delta v=\lambda/(2NT)$, where $\lambda$ is the wavelength, $N$ is the number of coherently processed pulses, and $T$ is the pulse repetition interval. Finally, the half-power angular resolution is defined as $\Delta\theta=0.886\lambda/D_{A}$, where $D_{A}$ is the size of the array’s aperture. For tracking, range, Doppler, and AoA are needed to determine target position and motion. For classification, additional discriminative features such as micro-Doppler signatures become critical for distinguishing between different target types or activities.
@@ II-D1 ISAC Performance Metrics
The performance of ISAC designs can be evaluated by characterizing fundamental trade-offs between sensing and communication metrics [5, 55, 56]. Generally, an ISAC metric combines communication and radar matrices from Tables III and IV, along with newly defined notions.
The concept of radar estimation information rate, introduced in [55, 57], parallels the data information rate in communication systems. It treats radar illumination on a target as a form of non-cooperative communication, where parameters such as delay, Doppler, and spatial directions are transmitted. Consequently, the radar channel can be viewed as a non-cooperative communication channel. This information rate is defined as the mutual information between radar and target, derived from the entropy of random estimation parameters and the entropy of estimation uncertainty [55]. Using the data processing inequality, the mutual information between the observation $Y$ and the true parameter $\theta$ is lower-bounded as [29] $I(Y;\theta)\geq\frac{1}{2}\mathrm{log}\left(\frac{P}{D}\right),$ where $\theta$ is Gaussian-distributed with variance $P$, and $D$ is the MSE distortion of estimating $\theta$. This bound links the equivalent estimation information rate to the MSE distortion.
@@ III Multi-Objective Optimization: Traditional versus Learning-Based Approaches
ISAC problems are inherently multi-objective optimization problems. Such a problem involves simultaneously optimizing two or more conflicting objectives which can be formulated as [58]
$\underset{\bm{x}}{\text{Minimize}}$ $\quad\bm{f}(\bm{x})=[f_{1}(\bm{x}),f_{2}(\bm{x}),\ldots,f_{i}(\bm{x})]$ (1) Subject to $\quad{g}_{j}(\bm{x})=0,\quad j=1,2,\ldots,m$ $\quad{h}_{k}(\bm{x})\leq 0,\quad k=1,2,\ldots,n\qquad\qquad$
in which $f_{i}(\bm{x})$ are the objective functions, $g_{j}(\bm{x})$ and $h_{k}(\bm{x})$ are the equality and inequality constraints, respectively, and $\bm{x}$ is the decision variable vector.
Unlike in single-objective problems, multi-objective optimization typically has no single global optimum. Instead, the goal is to identify a set of solutions that satisfy a chosen optimality criterion. The most common is Pareto optimality. A solution is Pareto-optimal if no other solution can improve one objective without degrading at least one other objective. There are several established approaches to solving multi-objective optimization problems [58, 59]. In this section, we discuss a few common methods, grouped into two categories: (i) traditional optimization methods, and (ii) learning-based techniques.
III-A Traditional Optimization Methods
For a given problem, the Pareto optimal set may contain infinitely many points. It is therefore important to distinguish between methods that generate the entire set (or a subset of it) and those that aim to obtain a single final solution. Using the scalarization method [59], we can transform a multi-objective optimization problem into a single-objective problem and find a single solution. This can be done in various ways [58]. We discuss weighted sum method and $\epsilon$-constraint method, as two prominent examples of vectorization. Other scalarization methods, such as the min–max method, lexicographic method, weighted product method, and goal programming exist and can be found in [58, 60].